OpenStudy (anonymous):

Please explain how to simplify: sqrt(16c^8)-sqrt(c^2)+2*sqrt(c^2)+sqrt(4c^2).

OpenStudy (luigi0210):

Remember that: \[\LARGE \sqrt{x}*\sqrt{x}=x\]

OpenStudy (luigi0210):

So \[\LARGE \sqrt{c^2}*\sqrt{c^2}=c^2\]

OpenStudy (anonymous):

that makes sense, like sqrt(25)*sqrt(25)=5?

OpenStudy (anonymous):

But how do I add and subtract them?

OpenStudy (jdoe0001):

\(\large \bf \sqrt{16c^8}-\sqrt{c^2}\times \sqrt{c^2}+\sqrt{4c^2}\quad ?\)

OpenStudy (luigi0210):

no, that would equal 25. Well, once you simplify them, if they have the same thing inside, then you can perform the indicated operation. So say we have: \[\LARGE 3\sqrt{x}+5\sqrt{x}\] We can add them because they both have that x. So it would just be \[\LARGE 8\sqrt{x}\]

OpenStudy (anonymous):

Close but not quite, \[\sqrt{16c^8}-\sqrt{c^2}+2\sqrt{c^2}+\sqrt4c^2\]

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

That makes sense.

OpenStudy (luigi0210):

So simplify the remaining two.

OpenStudy (anonymous):

Ok, just a sec.

OpenStudy (anonymous):

Sorry, I'm a little confused which two do you want me to simplify?

OpenStudy (luigi0210):

The \[\LARGE \sqrt{16c^8}\] and \[\LARGE \sqrt{4c^2}\]

OpenStudy (anonymous):

Oh, ok.

OpenStudy (anonymous):

\[\sqrt{20c^10}\]

OpenStudy (anonymous):

Ummm That one zero thing should be a ten.

OpenStudy (luigi0210):

Not exactly.

OpenStudy (anonymous):

Oh.

OpenStudy (anonymous):

So what did I do wrong?

OpenStudy (anonymous):

Hello?

OpenStudy (luigi0210):

Sorry about that, but here: \[\LARGE \sqrt{4*4c^8}-c^2+\sqrt{2*2c^2}\] Simplified: \[\LARGE 4c^4-c^2+2c\]

OpenStudy (anonymous):

Hmm, Okay, I think that makes sense. However, the third term should have been \[2\sqrt{c^2}\] not \[c^2\], or maybe I'm just confusing my self, lol.

OpenStudy (luigi0210):

No, it's c, because if we break it up: \[\LARGE \sqrt{c*c}\] and you know what that means?

OpenStudy (luigi0210):

\[\LARGE \sqrt{c}*\sqrt{c}\] is another way of writing it.

OpenStudy (anonymous):

No I didn't understand that, but I think that makes sense.

OpenStudy (anonymous):

hmm, ok now I get it.

OpenStudy (anonymous):

Sorry I'm not very quick at this.

OpenStudy (luigi0210):

It's alright.

OpenStudy (anonymous):

So, if I want to simplify sqrt(4c^2) it would look like 2*2c^2?

OpenStudy (luigi0210):

Breaking apart you mean?

OpenStudy (luigi0210):

*it apart

OpenStudy (anonymous):

Yes.

OpenStudy (luigi0210):

Well if you want to go down to the simplest: \[\LARGE \sqrt{2*2*c*c}\]

OpenStudy (anonymous):

Ok, that makes sense. Now what?

OpenStudy (luigi0210):

If there's 2 of the same things, you can take them out as one, so the 2 and c's are the same: \[\LARGE \sqrt{\color{red}{2*2}*\color{green}{c*c}}=\color{red}{2}*\color{green}{c}\]

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

So I have:\[4c^4-c^2+2c\] now?

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

I meant:

OpenStudy (anonymous):

\[4c^4-c^2+2c+2c\]

OpenStudy (luigi0210):

Well, since none of them have the same exponents, you just leave it as it is. So \[\LARGE 4c^4-c^2+2c\]Will be your answer

OpenStudy (anonymous):

Because I can't combine like terms, right?

OpenStudy (luigi0210):

Exactly.

OpenStudy (anonymous):

Great, thank you so much for your help. :)

OpenStudy (anonymous):

I understand a lot better now.

OpenStudy (luigi0210):

Anytime bud~